OpenStudy (anonymous):

Just need help with this problem (x^-3)^-5x^6

OpenStudy (anonymous):

\[ (x^{-3})^{-5}x^6=(x^{-3 \times( -5)})x^6 =(x^{15})x^6 =x^{15+6} =x^{21}\]

OpenStudy (anonymous):

@stanstan

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@stanstan medal please